pandasurvive:

A flare was launched straight up from the ground with an initial velocity of 176 ft/s and returned to the ground after 11 s. The height of the flare t seconds after launch is modeled by the function f(t)=−16t^2+176t . What is the maximum height of the flare, in feet?

AngeI:

@tigerlover

tigerlover:

One way to figure out the maximum height is to take the derivative of the equation. So, f'(t)=-32t+176 Find for what values of t f'(t) is 0, as that is the maximum/minimum of your function (when the slope of the tangent is 0). So 0=-32t+176 -32t=-176 t=5.5 s. So at 5.5 s is when a max height takes place. Plug it back into your original equation, and you should get: f(5.5)=-16(5.5)^2+176(5.5) f(5.5)=484 ft

tigerlover:

Ahhhh

tigerlover:

WHo dat?